A turntable of mass 4.90 kg has a radius of 0.079 m and spins with a frequency of 0.532 rev/s. What is its angular momentum? Assume the turntable is a uniform disk. Incorrect: Your answer is incorrect. kg
Answer #1

Given

Mass , m = 4.90 kg

Radius , r = 0.079 m

frequency , f = 0.532 rev /s

Angular velocity , ? = 2? (0.532 ) = 3.343 rad/s

Moment of inertia of uniform disk, I = 1/2 mr^2

                                                          = 0.5 *4.9 kg * (0.079 m)^2

                                                       I = 0.0153 kg m^2

Angular momentum, L = I ?

                                   = 0.0153 kg m^2 * 3.343 rad/s

                                 L = 0.0511 kg m^2 / s
Answer #2

  ? = m/?r^2
dI = ?r^2drd?
I = (2/3)??r^3 = (2/3)mr
P = I? = (2/3)mr?
P = 2*4.90*0.079*0.532/3
P = 0.1373 kg-m/s
Answer #4

Given m=4.90kg, r=0.079m, ?=0.532 rev/s
For a thin, solid disk:
I = mr